13. From what height above the ground was it released? a. 12 m b. 18 m c. 24 m A mass is released from rest and strikes the ground with an impact speed of 25 m/s. d. 32 m e. 40 m
Answer #1

Here we apply conservation of energy equation

That implies, Potential energy, Ui = Kinetic energy KEf

Therefore, m*g*h = 0.5*m*v²

So, height, h = 312.5/9.81 = 32 m (approximately)